# Renormalizable Spectral-Shell Dynamics as the Origin of Neural Scaling Laws

### v4

@@ 8 Conclusion
Appendix A Exact Shell-Energy Bookkeeping from Modewise Dynamics
This appendix derives the exact shell-level energy balance underlying the main text. Its purpose is purely structural: to show how the quadratic shell energies used throughout the paper arise rigorously from the exact modewise dynamics. No weak-coupling, locality, or continuum assumptions are imposed here. All scaling-law results rely only on the renormalizable shell-dynamics assumption stated in Definition 4.1.
A.1 Modewise dynamics and antisymmetric coupling
Recall from Section 3.1 that, in the instantaneous eigenbasis of $M(t)$, the error amplitudes satisfy the exact system
$\partial_{t}g_{u}(t)+\sum_{v\neq u}g_{v}(t)\,\Omega_{v\to u}(t)=-\lambda_{u}(t)\,g_{u}(t),$ (16)
with coupling coefficients
$\Omega_{v\to u}(t)=\frac{\langle\varphi_{u}(t),\dot{M}(t)\varphi_{v}(t)\rangle}{\lambda_{v}(t)-\lambda_{u}(t)},\qquad u\neq v.$
Since $\dot{M}(t)$ is self-adjoint, these coefficients satisfy the strict antisymmetry
$\Omega_{v\to u}(t)=-\Omega_{u\to v}(t),\qquad u\neq v.$ (17)
A.2 Logarithmic shells and quadratic energies
Fix a logarithmic partition of the spectrum as in Section 3.2:
$S_{\alpha}:=\{u:\ \lambda_{u}(t)\in[\lambda_{\alpha},\lambda_{\alpha+1})\},\qquad\lambda_{\alpha}=\lambda_{0}q^{\alpha},\quad q>1.$
For each shell, define the quadratic shell energy
$E_{\alpha}(t):=\frac{1}{2}\sum_{u\in S_{\alpha}}g_{u}(t)^{2}.$ (18)
This choice is canonical, since the total training loss is exactly
$\mathcal{L}(t)=\frac{1}{2}\|e_{t}\|_{2}^{2}=\sum_{\alpha}E_{\alpha}(t).$
A.3 Exact cancellation of shell-internal coupling
Differentiate (18):
Substituting (16) yields
$\frac{d}{dt}E_{\alpha}(t)=-\sum_{u\in S_{\alpha}}\lambda_{u}(t)\,g_{u}(t)^{2}-\sum_{u\in S_{\alpha}}\sum_{v\neq u}g_{u}(t)g_{v}(t)\,\Omega_{v\to u}(t).$ (19)
Split the coupling sum into contributions with $v\in S_{\alpha}$ and $v\notin S_{\alpha}$. By antisymmetry (17), the shell-internal contribution vanishes identically:
$\sum_{u\in S_{\alpha}}\sum_{\begin{subarray}{c}v\in S_{\alpha}\\ v\neq u\end{subarray}}g_{u}g_{v}\,\Omega_{v\to u}=0.$ (20)
Thus shell-internal coupling acts as a pure rotation of amplitudes and does not change the quadratic shell energy.
A.4 Exact shell-energy balance
Only dissipation and inter-shell exchange remain. Define the inter-shell quadratic-energy flux
$\mathcal{F}_{\beta\to\alpha}(t):=-\sum_{\begin{subarray}{c}u\in S_{\alpha}\\ v\in S_{\beta}\end{subarray}}g_{u}(t)g_{v}(t)\,\Omega_{v\to u}(t).$
Then (19) reduces to the exact identity
Moreover, antisymmetry implies the action–reaction property
$\mathcal{F}_{\beta\to\alpha}(t)=-\mathcal{F}_{\alpha\to\beta}(t),$
so that inter-shell coupling conserves total quadratic energy:
$\sum_{\alpha}\sum_{\beta\neq\alpha}\mathcal{F}_{\beta\to\alpha}(t)=0.$
The only mechanism that decreases $\sum_{\alpha}E_{\alpha}(t)$ is dissipation.
A.5 Relation to renormalizable shell dynamics
Equation (21) is an exact bookkeeping identity. It involves no approximation, no continuum limit, and no assumption on the structure of the coupling coefficients beyond antisymmetry.
The renormalizable shell-dynamics assumption introduced in Definition 4.1 is a macroscopic closure of this identity: it postulates that, after coarse-graining, the cumulative effect of the inter-shell fluxes can be summarized by a controlled net boundary flux plus a negligible remainder. All scaling-law results in the main text are derived from this shell-level closure and do not rely on any further microscopic characterization.
Appendix B Subcritical Drift Regimes ($b\leq 1$)
In the main text, we focus on the supercritical regime $b>1$, under which the spectral drift induces a progressive resolution cascade and gives rise to well-defined data scaling laws. For completeness, we briefly discuss the remaining cases $b\leq 1$ here, which correspond to qualitatively different learning dynamics.
Critical regime ($b=1$).
When $b=1$, the characteristic equation
$\frac{d\lambda}{dt}=-c(t)\lambda$
admits an exponential solution $\lambda(t)=\lambda_{0}e^{-\tau(t)}$, where $\tau(t)=\int_{0}^{t}c(s)\,ds$. In this regime, all spectral modes decay at the same relative rate, and no resolution hierarchy is formed. This behavior corresponds to the well-known lazy or NTK regime, in which the induced kernel remains effectively fixed and feature learning is absent. While learning still occurs, it does not exhibit progressive resolution unlocking or nontrivial scaling with data.
Subcritical regime ($0<b<1$).
For $0<b<1$, the characteristic flow
$\lambda(t)^{1-b}=\lambda_{0}^{1-b}-(1-b)\tau(t)$
reaches the spectral boundary $\lambda=0$ in finite effective time. In practical settings, this boundary is replaced by a resolution floor determined by finite model capacity, numerical precision, or noise. As a result, the spectral mass is rapidly transported toward this floor, after which no further refinement of resolution is possible.
Importantly, learning is not absent in this regime. Rather, the model converges quickly to a fixed representational limit, beyond which increasing the amount of data only accelerates convergence toward the same solution. In particular, no asymptotic scaling law with respect to data size can be sustained. In this sense, the learning dynamics in the subcritical regime are not scalable.
Implications.
The regimes $b\leq 1$ thus correspond to either non-hierarchical (lazy) or non-scalable learning dynamics. Since modern deep learning systems of interest empirically exhibit strong scalability with data, we restrict our main analysis to the supercritical regime $b>1$, and defer the discussion of subcritical cases to this appendix.
Appendix C The polynomial bound of effective time $\tau(t)$
Theorem 2 (Polynomial Growth of the Effective Time under Standard Optimization Assumptions).
Let the drift strength $c(t)$ of the spectral velocity $v(\lambda,t)\sim-c(t)\lambda^{b}$ be determined by the gradient statistics of a neural network trained with a stable optimizer (SGD, AdamW, momentum SGD) under bounded gradient norms and non-exploding learning-rate schedules. Then the effective time
$\tau(t)=\int_{0}^{t}c(s)\,ds$
satisfies
$\tau(t)=t^{\alpha}\,L(t),\qquad\alpha>0,$
where $L(t)$ is a slowly varying function in the sense of Karamata:
$\lim_{t\to\infty}\frac{L(kt)}{L(t)}=1\qquad\forall k>0.$
In particular, $\tau(t)$ is asymptotically polynomial up to a slowly varying correction.
Proof.
The drift amplitude $c(t)$ is determined by
$v(\lambda,t)=-c(t)\lambda^{b},\qquad c(t)=\langle\phi_{\lambda}(t),\dot{M}(t)\,\phi_{\lambda}(t)\rangle\,\lambda^{-b}.$
The operator derivative $\dot{M}(t)$ equals the expected gradient outer product
$\dot{M}(t)=\mathbb{E}\bigl[\nabla_{\theta}f(x;\theta(t))\,\nabla_{\theta}f(x;\theta(t))^{\top}\bigr]\cdot\eta(t),$
where $\eta(t)$ is the learning rate.
Under standard stability conditions of deep learning:
  • bounded gradient norms: $\|\nabla_{\theta}f(x;\theta(t))\|\leq C$,
  • non-exploding learning rates,
  • Lipschitz continuity of the gradient field,
  • noise-controlled SGD (bounded variance),
the quantity $c(t)$ is positive and satisfies
$c_{\min}\,t^{\alpha-1}\leq c(t)\leq c_{\max}\,t^{\alpha-1}L(t),$
for some $\alpha>0$ and a slowly varying function $L$. This follows from classical results on stochastic approximation and ODE–SDE averaging, which imply that gradient magnitudes evolve according to polynomially bounded envelopes and admit slowly varying corrections through learning-rate modulation and noise.
Integrating $c(t)$, we obtain
$\tau(t)=\int_{0}^{t}c(s)\,ds=t^{\alpha}\,L(t),$
where the same $L(t)$ (up to asymptotic equivalence) is slowly varying because integration preserves slow variation.
Thus the effective time is asymptotically a polynomial multiplied by a slowly varying factor. ∎

### v5

@@ 9 Conclusion
Appendix A KL-to-MSE justification for log-probability error
This appendix justifies the quadratic error observable used in the empirical spectral-shell analysis. In language modeling with one-hot supervision, the ground-truth conditional distribution is concentrated at the correct token $y^{*}$, while the model predicts $q_{\theta}(\cdot\mid x)$. For a fixed context $x$, the token-level KL divergence satisfies
$\mathrm{KL}\!\big(p(\cdot\mid x)\,\|\,q_{\theta}(\cdot\mid x)\big)=\sum_{y}p(y\mid x)\log\frac{p(y\mid x)}{q_{\theta}(y\mid x)}=-\log q_{\theta}(y^{*}\mid x)+C(x),$
where $C(x)$ depends only on $p(\cdot\mid x)$ and is independent of $\theta$.
Define the (log-probability) error observable
$e(x)\;:=\;-\log q_{\theta}(y^{*}\mid x).$
When the model is already assigning high probability to the correct token, $q_{\theta}(y^{*}\mid x)\approx 1$, we have $e(x)\approx 0$ and a local quadratic approximation is appropriate. Consider $q=\exp(-e)$, so that $q\approx 1$ iff $e\approx 0$. Expanding $\exp(-e)$ around $e=0$ gives
$q=e^{-e}=1-e+\tfrac{1}{2}e^{2}+O(e^{3}),\qquad\Rightarrow\qquad 1-q=e+O(e^{2}).$
Thus, in a neighborhood of $q=1$, any smooth loss functional that is minimized at $q=1$ admits a second-order expansion in $e$. In particular, the KL divergence above equals $e$ up to an additive constant, and its local quadratic surrogate takes the form
$\mathrm{KL}(p\|q_{\theta})\;\approx\;\tfrac{1}{2}\,e(x)^{2}\qquad\text{for }e(x)\ll 1\;(q_{\theta}(y^{*}\mid x)\approx 1).$
Therefore, using the squared log-probability $e(x)^{2}$ as a mean-squared error observable is consistent with the quadratic-loss framework that underlies the spectral-shell bookkeeping: the induced operator dynamics and shell-energy definitions apply directly once the error signal is taken to be $e(x)$ and the energy density is built from $e(x)^{2}$.
Remark (what is measured empirically).
All empirical spectra in this paper are computed from the squared error $e(x)^{2}=\big(-\log q_{\theta}(y^{*}\mid x)\big)^{2}$ and then aggregated into logarithmic spectral shells.
Appendix B Spectral estimation via (stochastic) Lanczos and shell binning
This appendix describes the Lanczos-based estimator used to obtain log-shell energies without forming the operator $M=J_{\theta}J_{\theta}^{*}$ explicitly.
B.1 Goal: shell energies from the spectral measure of $M$
Fix a checkpoint $\theta$ and let $M:=J_{\theta}J_{\theta}^{*}$ be the PSD operator governing error dynamics in function space. Let $e$ denote the error signal (Section LABEL:sec:setup; in our experiments $e(x)=-\log q_{\theta}(y^{*}\mid x)$). Formally diagonalizing $M\phi_{u}=\lambda_{u}\phi_{u}$ and expanding $e=\sum_{u}g_{u}\phi_{u}$, the shell-integrated energy is
$E_{\alpha}\;:=\;\sum_{u:\,\lambda_{u}\in[\lambda_{\alpha},\lambda_{\alpha+1})}g_{u}^{2}.$
Equivalently, define the spectral measure $\mu_{e}$ of $M$ induced by $e$ so that for any test function $f$,
$\langle e,f(M)e\rangle=\int f(\lambda)\,d\mu_{e}(\lambda),\qquad d\mu_{e}(\lambda)=\sum_{u}g_{u}^{2}\,\delta(\lambda-\lambda_{u})\,d\lambda.$
Then $E_{\alpha}$ is exactly the mass of $\mu_{e}$ in the bin $\lambda_{\alpha},\lambda_{\alpha+1})$.
B.2 Lanczos tridiagonalization and quadrature
Lanczos constructs, from repeated applications of $M$ to vectors, an orthonormal basis of the Krylov subspace $\mathcal{K}_{m}(M,v_{0})=\mathrm{span}\{v_{0},Mv_{0},\dots,M^{m-1}v_{0}\}$ and a symmetric tridiagonal matrix $T_{m}\in\mathbb{R}^{m\times m}$ such that
$MV_{m}\;\approx\;V_{m}T_{m},\qquad V_{m}=[v_{0},\dots,v_{m-1}],\quad V_{m}^{*}V_{m}=I.$
We choose the starting vector to be the normalized error direction
$v_{0}:=\frac{e}{\|e\|}.$
Let $T_{m}=U\,\mathrm{diag}(\tilde{\lambda}_{i})\,U^{*}$ be the eigendecomposition of the tridiagonal matrix. Then Gauss quadrature yields an approximation of the spectral measure $\mu_{e}$ by the discrete measure
$\mu_{e}(d\lambda)\;\approx\;\|e\|^{2}\sum_{i=1}^{m}w_{i}\,\delta(\lambda-\tilde{\lambda}_{i})\,d\lambda,\qquad w_{i}:=(U_{1i})^{2},$
where $U_{1i}$ is the first component of the $i$-th eigenvector of $T_{m}$.
B.3 Log-shell binning
Given logarithmic bin edges $\{\lambda_{\alpha}\}_{\alpha=0}^{B}$, define the estimated shell energies by accumulating quadrature masses:
$\widehat{E}_{\alpha}\;:=\;\|e\|^{2}\sum_{i:\,\tilde{\lambda}_{i}\in[\lambda_{\alpha},\lambda_{\alpha+1})}w_{i}.$
We report $\log\widehat{E}_{\alpha}$ as a function of $\log\lambda_{\alpha}^{\mathrm{ctr}}$, where $\lambda_{\alpha}^{\mathrm{ctr}}$ is the geometric bin center.
B.4 Data-induced averaging (no extra randomness)
We do not rely on a global random seed. Variance reduction is achieved by averaging the resulting shell estimates across many independent data batches: each batch provides an empirical error vector $e$ (computed from a large set of tokens), and the above Lanczos+binning pipeline produces $\widehat{E}_{\alpha}$ for that batch; we then average $\widehat{E}_{\alpha}$ over batches.
B.5 Pseudo-code
Algorithm 1 Lanczos shell-energy estimation at a fixed checkpoint
1:Operator-vector product routine $u\mapsto Mu$; error vector $e$; Lanczos steps $m$; log-shell edges $\{\lambda_{\alpha}\}_{\alpha=0}^{B}$.
2:Estimated shell energies $\widehat{E}_{\alpha}$.
[…]
4:for $k=0,\dots,m-1$ do
5:  $w\leftarrow Mv_{k}-\beta_{k}v_{k-1}$
6:  $\alpha_{k}\leftarrow\langle v_{k},w\rangle$
7:  $w\leftarrow w-\alpha_{k}v_{k}$
8:  $\beta_{k+1}\leftarrow\|w\|$
9:  if $\beta_{k+1}=0$ then break
10:  end if
11:  $v_{k+1}\leftarrow w/\beta_{k+1}$
12:end for
13:Form tridiagonal $T$ with diagonal entries $\alpha_{k}$ and off-diagonals $\beta_{k+1}$
14:Compute eigendecomposition $T=U\,\mathrm{diag}(\tilde{\lambda}_{i})\,U^{*}$
15:Set weights $w_{i}\leftarrow(U_{1i})^{2}$
16:Initialize $\widehat{E}_{\alpha}\leftarrow 0$ for all $\alpha=0,\dots,B-1$
17:for $i=1,\dots,m$ do
18:  Find $\alpha$ such that $\tilde{\lambda}_{i}\in[\lambda_{\alpha},\lambda_{\alpha+1})$
19:  if such $\alpha$ exists then
20:   $\widehat{E}_{\alpha}\leftarrow\widehat{E}_{\alpha}+\|e\|^{2}\,w_{i}$
21:  end if
22:end for
23:return $\{\widehat{E}_{\alpha}\}_{\alpha=0}^{B-1}$
Implementation note.
The only required primitive is the product $u\mapsto Mu=J_{\theta}(J_{\theta}^{*}u)$. In modern autodiff frameworks this can be implemented via a VJP followed by a JVP, without constructing $J_{\theta}$ or $M$ explicitly.